OpenStudy (anonymous):

What is the solution of the system of y=x - 4 and y=3x

OpenStudy (anonymous):

Plug in 3x into the equation y = x -4, so it would be: 3x = x - 4 Then all you have to do is solve; 3x = x -4 -x -x 2x = -4 [x = -2] Then plug the x value in the equation again... y = 3 x y = 3 * -2 [y = -6] (Answers are in []s)